OpenStudy (anonymous):

14=67y+78x how'd i solve for x? PLEASE HELP!!!!!!!!

OpenStudy (anonymous):

do i subtract 78x?

OpenStudy (anonymous):

Help Please!@susiegeiger1113 and @brandi2013

OpenStudy (anonymous):

and @Gracie112 ?!?

OpenStudy (anonymous):

I worked it out, and it doesn't make any sense to me. So I'm going to find X this time around so hang on a sec.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Just do it like this, \[x=y+1\]Equate it to x, and the rest move it to the other side.

OpenStudy (anonymous):

Okay so solving for X right away didn't work either, but what I've been trying to make work is subtracting Y or X to get the other..so like 67y = 14 -78x, dividing by 67. X = 14/78 - 67/78y. Fill that into the original equation, 67y + 78(14/78 - 67/78y) = 14. 67y + 14 - 67 = 14

OpenStudy (anonymous):

*67y + 14 - 67y = 14. 0y = 0..y = 0. Fill back in to original, X = 14/78 or 7/39

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[67y+78x=14\]From my basic approach, move everything to the right EXCEPT x. Then we got...\[78x=14-67y\] Then, make x=1 just by dividing the LEFT and the RIGHT by 78. that way you'll get x

OpenStudy (anonymous):

? i confused!

OpenStudy (anonymous):

If you want just X, you would have to do things my way but if you want to still have an equation with y in it, you would do Lynncake's way. Her way would give you x = 7/39 - 67/78y

OpenStudy (anonymous):

okay.. i get it.. i need it how you do it @brandi2013 but thanks @Lynncake